OpenStudy (anonymous):

Pyramid A is a square pyramid with a base side length of 12 inches and a height of 8 inches. Pyramid B is a square pyramid with a base side length of 36 inches and a height of 24 inches. How many times bigger is the volume of pyramid B than pyramid A?

OpenStudy (tkhunny):

Perhaps you could calculate the two volumes and see?

OpenStudy (tkhunny):

On the other hand: Volume A = \(\dfrac{1}{3}\cdot Height\cdot Side^{2}\) 36/12 = 3 24/8 = 3 Volume B = \(\dfrac{1}{3}\cdot (3\cdot Height)\cdot (3\cdot Side)^{2} = 9\;\cdot\) Volume A THAT was cool. Think on that.

OpenStudy (anonymous):

Yeah, i fiugred that out, thank you :)

OpenStudy (anonymous):

do you think you can help with another question? i am having a difficult time. @tkhunny

OpenStudy (anonymous):

An hourglass consists of two sets of congruent composite figures on either end. Each composite figure is made up of a cone and a cylinder, as shown below: Hourglass with sand measuring 47 millimeters high © 2011 Jupiterimages Corporation Each cone of the hourglass has a height of 12 millimeters. The total height of the sand within the top portion of the hourglass is 47 millimeters. The radius of both the cylinder and cone is 4 millimeters. Sand drips from the top of the hourglass to the bottom at a rate of 10π cubic millimeters per second. How many seconds will it take until all of the sand has dripped to the bottom of the hourglass? 6.4 62.4 8.5 56.0

OpenStudy (tkhunny):

You may have figured it out. Too bad you shared nothing with us. Generally, one question per post. I got bored reading it. You'll have to provide a diagram and then SHOW YOUR WORK. No excuses.